OpenStudy (anonymous):

use the one to one property tp solve log_2(x-3)-log_2(2x+1)=-log_2(4).. the answer in the book is 13/2.. but i dont know how to get there so please show work!! thank you!!

OpenStudy (anonymous):

@phi .. you helped me last time and was very helpful.. any ideas on thsi one?

OpenStudy (phi):

on the right side log_2(4) means what number x is it so that 2^x = 4 any idea ?

OpenStudy (anonymous):

2

OpenStudy (phi):

\[ \log_2(x-3)-\log_2(2x+1)=-\log_2(4) \\ \log_2(x-3)-\log_2(2x+1)= -2 \]

OpenStudy (phi):

On the left side, use log(a) - log(b) = log(a/b) to combine the two logs

OpenStudy (anonymous):

okay ya i did that!

OpenStudy (anonymous):

i got... log_2((x-3)/(2x+1))=-2 .. can i just cross out that log_2?

OpenStudy (anonymous):

@phi

OpenStudy (phi):

to get rid of a log, you make it the exponent of the base. in other words, do this on both sides: \[ 2^{\log_2(stuff)} = 2^{-2} \] by definition, the left side turns into \[ stuff = 2^{-2} \]

OpenStudy (anonymous):

wait sorry im really confused at that point

OpenStudy (raden):

if log_2((x-3)/(2x+1))=-2 then (x-3)/(2x+1)=2^-2 solve for x

OpenStudy (anonymous):

so on the right itd be .25?

OpenStudy (phi):

yes, this part is confusing. on the right side you get ¼ (or 0.25) but for the time being I would use ¼ the left side is confusing, but the idea is if you have log_2(A) you can "undo" the log_2 by making log_2(A) an exponent of 2 \[ 2^{\log_2(A)} \] which *by definition* is A

OpenStudy (phi):

so we make each side the exponent of 2 (to keep things equal) the right side turns into ¼ the left side turns into (x-3)/(2x+1)

OpenStudy (anonymous):

ok i got thattt now just solve?

OpenStudy (anonymous):

wouldn't i just multiply both sides by (2x+1) next? so itd cancel out on the left and then distrbute on the right?

OpenStudy (phi):

yes, sounds like a good plan and multiply both sides by 4 to get rid of the fraction

OpenStudy (anonymous):

what one should i do first?

OpenStudy (phi):

order does not matter. you can think of it as cross multiplying (if you like)

OpenStudy (phi):

\[ \frac{(x-3)}{(2x+1)}=\frac{1}{4} \]

OpenStudy (anonymous):

so on the right id just have 1.. then on the left id have,, (4x-12)/ 8x+8)?

OpenStudy (anonymous):

ohh wait so i can just corss multiply that equation up top?

OpenStudy (phi):

no, you multiply both sides by 4. the 4 is up top.

OpenStudy (phi):

**no to (4x-12)/ 8x+8)? yes to cross multiply.

OpenStudy (anonymous):

so.... (4x-12)/(2x+1)

OpenStudy (phi):

(4x-12)/(2x+1)=1 always write the whole equation

OpenStudy (phi):

now multiply both sides by (2x+1)

OpenStudy (anonymous):

so 8x^2-12=2x+1... right.. or do i foil

OpenStudy (phi):

notice that (2x+1) times the left side will cancel out the (2x+1) on the bottom

OpenStudy (phi):

you could multiply (2x+1)(4x-12) and use FOIL (which gives a different answer than what you got), but that is not what we want. \[\cancel{ (2x+1)} \cdot \frac{4x-12}{\cancel{2x+1}} = (2x+1) \cdot 1 \]

OpenStudy (anonymous):

thank youu!!!!!!1!

OpenStudy (anonymous):

how about find the x intercepts of the function: f(x)=e^(x+4)-e... the answers in the book is (-3,0).. any idea how to perform it? @phi

OpenStudy (phi):

x intercepts are where y ( i.e. f(x) ) is 0 so we have to solve e^(x+4) - e^1 = 0 we could factor out e^1 from each term remember \( e^{a+b}= e^a \cdot e^b \) lets write x+4 as (x+3) + 1 so we have \[ e^{(x+3)+1} - e^1 =0\\ e^{(x+3)}\cdot e^1-e^1 = 0 \]

OpenStudy (anonymous):

ok got that!

OpenStudy (anonymous):

what would i do nextrt

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

so then just (x+3)=0 and it would be -3?

OpenStudy (phi):

starting with \[ e^{(x+3)}\cdot e^1-e^1 = 0 \] what do you get after factoring out e^1 ?

OpenStudy (anonymous):

ummm e^(x+3)=0? wwouldnt both of the e^1 factor out?

OpenStudy (phi):

what would you get if you factor a out of (ba - a) = 0

OpenStudy (anonymous):

a(b-1)=0

OpenStudy (phi):

now notice a could be e^1 and b could be e^(x+3)

OpenStudy (phi):

it is the same problem, with different symbols.

OpenStudy (anonymous):

soo.. e^1(e^(x+3)-1)=0

OpenStudy (phi):

yes. now divide both sides by e^1 (or multiply both sides by e^(-1) )

OpenStudy (anonymous):

okk

OpenStudy (phi):

or in a(b-1)= 0 divide both sides by a

OpenStudy (anonymous):

so.. (e^(x+3)-1)=e^1

OpenStudy (phi):

0/e^1 is not e^1

OpenStudy (anonymous):

0

OpenStudy (phi):

\[ \frac{1}{\cancel{e^1}}\cdot \cancel{e^1}(e^{(x+3)}-1)=\frac{0}{e^1} \]

OpenStudy (phi):

you now have \[ e^{x+3}-1=0 \] add 1 to both sides

OpenStudy (anonymous):

ok

OpenStudy (phi):

we have so far ?

OpenStudy (anonymous):

e^(x+3)=1

OpenStudy (anonymous):

@phi

OpenStudy (phi):

any idea how to "bring down" the (x+3) ? I would take the natural log of both sides.

OpenStudy (anonymous):

yaa i was gonna say that.. but then what happens to the ln1 on the right? how to i cancel the ln out

OpenStudy (phi):

you don't cancel out ln 1 however, we can figure out what number that is. what x is it , so that e^x = 1 ?

OpenStudy (anonymous):

e^0 =1

OpenStudy (anonymous):

sweet thanks i got it!!

OpenStudy (phi):

and ln(e^0) = ln 1 or 0 * ln(e) = ln(1) 0 = ln(1)